"5 years ago mother was 7 times as old as her daughter. 5 years hence she will be 3 times as old as her daughter. find their present age?"
Answer by Aswathy(23)

Let daughter's and mother's present ages be x and y respectively.
As,5 years ago mother's age was 7 times her daughter's age, so,mother's will be y-5=7(x-5)
Similarly,5 years from now the mother's age will be y+5=3(x+5)
By solving both the equations we get x=10 & y=40.
So the present ages of mother and her daughter are 40 & 10 years respectively.